OpenStudy (anonymous):

Use the quadratic formula to find the roots of the equation. Round to tenths if necessary. x2 - 6x + 2 = 0 A.{0.4, 5.6} B.{0.5, 4.8} C.{1.4, 2.4} D.no real number solution

OpenStudy (anonymous):

someone please help me with this i'm very confused

OpenStudy (anonymous):

hint... use this formula: \[\LARGE {x_{1/2}} = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\] and you have..a=1 b=-6 c=2

OpenStudy (anonymous):

\[\LARGE \Delta = {b^2} - 4ac\] \[\LARGE \begin{array}{l}\Delta > 0\;\;\;\;{\rm{ two solutions}}{\rm{.}}\\\Delta {\rm{ = 0\;\;\;\; one solution}}{\rm{.}}\\\Delta {\rm{ < 0 \;\;\;\; no solution}}{\rm{.}}\end{array}\]

OpenStudy (anonymous):

 that's what your saying?